OpenStudy (anonymous):

please help , limit n tends to infinity log n/n^0.1 =?? ^=to the power

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty } (\log n)/n ^{0.1}\]

OpenStudy (cwrw238):

try l'hopitals rule

OpenStudy (anonymous):

but isn't it for log n/n here the power of n=0.1 how cn we apply L'hospital rule here ?

OpenStudy (cwrw238):

l'hopitals rule applies for function which are indeterminate like 0/0 amd inf / inf yea you are probably right - i studied limits last year in my maths course and i'm afraid they are not my strong point.

OpenStudy (anonymous):

ohh !!

OpenStudy (cwrw238):

sorry

OpenStudy (anonymous):

np, thank u :)

OpenStudy (cwrw238):

i know that limit of log n as n approaches infinity is infinity

OpenStudy (anonymous):

then i think we cn apply L'hospital rule as it is an indeterminatistic form !

OpenStudy (cwrw238):

ok lets try it differentiating = 1/n ------ 0.1n^-0.9 = 1 ------- o.1n * n ^-0.9 = 1 / 0.1 n^0.1 = 0 as n approaches infinity

OpenStudy (cwrw238):

I checked it out on wolframalpha - thats the right answer

OpenStudy (anonymous):

thank u :)